OpenStudy (anonymous):

solve... 7/x=8/x-1/11

OpenStudy (anonymous):

\[7/x - 8/x = -1/11\] \[-1/x = -1/11\] so x = 11

OpenStudy (anonymous):

i got -1

OpenStudy (anonymous):

plug in your answer into the original equation, -1 will not work. x = 11

OpenStudy (anonymous):

ty